# Sample, Don’t Search: Rethinking Test-Time Alignment for Language Models

### v1

@@ Appendix B Computational Cost of QAlign
As outlined in Faria et al. (2024), when using the suffix proposal distribution from Quest, each step samples, on average, an index at the midpoint of the sentence from the uniform distribution. Assuming a fixed sentence length $N$ for simplicity, this requires generating only $\frac{N}{2}$ new tokens on average per step. With typical transformer-based models, this allows us to reuse the majority of the key-value cache from previous iterations in both the base model and reward model. The reuse extends even beyond half of the computation due to the static prompt. However, unlike when generating independent samples, the computation for each prompt is sequential. We need to first generate the sentence $y^{t}$ before generating the sentence $y^{t+1}$. This means that regardless of the compute capability, compared with sampling independent samples, there is always an inherent latency overhead.
@@
In summary, for a chain of $T$ steps, QAlign is expected to generate $\frac{(T+1)N}{2}$ tokens in total: $N$ tokens for the initial hypothesis, plus an average of $\frac{N}{2}$ tokens for each of the remaining $T-1$ steps. In comparison, generating $T$ independent samples requires decoding $T\times N$ tokens. Therefore, for an equal number of samples, Quest requires approximately half as many tokens from generation alone than sampling independently, translating to roughly half the FLOPs.
Appendix C Empirical Observations on Distribution of Rewards
@@
When analyzing reward model predictions across independently generated responses from the base model for individual prompts, we consistently observed a bimodal distribution forming a two-component Normal mixture with distinct means and variances (Figure 5). This pattern appeared across all three reward models used.
@@
Figure 5: Histogram of rewards assigned by Tülu3-8B-RM to $1,024$ responses generated by Tülu3-8B-SFT for $9$ randomly sampled prompts from GSM8K. For each prompt, we fit a two-component Gaussian mixture model to characterize the reward distribution.
@@
Appendix D Prompts Used for Evaluation
@@ D.2 MATH-500 Dataset
D.3 Multiple Choice Datasets (TQA, MMLU)
@@
Appendix E Full General Alignment Plots
@@
Figure 6: Error rate across multiple evaluation datasets (GSM8K, MATH500, MMLU-Redux, TruthfulQA, and IFEval) as a function of the floating point operations (FLOPS) in log scale. We compare $\bullet$ QAlign with Tülu3-8B-SFT against four baselines: $\blacktriangle$ majority vote (MV) Tülu3-8B-DPO, and applied to Tülu3-8B-SFT the methods $\bullet$ best-of-$n$ (BoN), $\bullet$ MV, and $\bullet$ weighted MV (WMV) . All experiments use temperature 1.0 with reasoning included in model outputs. Note that Tülu3-8B-DPO model is the result of doing preference finetuning on the Tülu3-8B-SFT with 271k preference pairs. The costs associated with this process are not accounted for in this plot.

### v2

@@ Appendix D Computational Cost of QAlign
As outlined in quest, when using the suffix proposal distribution from Quest, each step samples, on average, an index at the midpoint of the sentence from the uniform distribution. Assuming a fixed sentence length $N$ for simplicity, this requires generating only $\frac{N}{2}$ new tokens on average per step. With typical transformer-based models, this allows us to reuse the majority of the key-value cache from previous iterations in both the base model and reward model. The reuse extends even beyond half of the computation due to the static prompt. However, unlike when generating independent samples, the computation for each prompt is sequential. We need to first generate the sentence $y^{t}$ before generating the sentence $y^{t+1}$. This means that regardless of the compute capability, compared with sampling independent samples, there is always an inherent latency overhead.
@@
In summary, for a chain of $T$ steps, QAlign is expected to generate $\frac{(T+1)N}{2}$ tokens in total: $N$ tokens for the initial hypothesis, plus an average of $\frac{N}{2}$ tokens for each of the remaining T1T-1 steps. In comparison, generating $T$ independent samples requires decoding $T\times N$ tokens. Therefore, for an equal number of samples, Quest requires approximately half as many tokens from generation alone than sampling independently, translating to roughly half the FLOPs.
Appendix E Empirical Observations on Distribution of Rewards
@@
When analyzing reward model predictions across independently generated responses from the base model for individual prompts, we consistently observed a bimodal distribution forming a two-component Normal mixture with distinct means and variances (Figure 5). This pattern appeared across the reward models used.
@@
Figure 5: Histogram of rewards assigned by Tülu3-8B-RM to $1,024$ responses generated by Tülu3-8B-SFT for 99 randomly sampled prompts from GSM8K. For each prompt, we fit a two-component Gaussian mixture model to characterize the reward distribution.
@@
Appendix F Prompts Used for Evaluation
@@ F.2 MATH-500 Dataset
F.3 Multiple Choice Datasets (TQA, MMLU)
@@
Appendix G Full General Alignment Plots
@@
Figure 6: Error rate across multiple evaluation datasets (GSM8K, MATH500, MMLU-Redux, TruthfulQA, and IFEval) as a function of the floating point operations (FLOPS) in log scale. We compare $\bullet$ QAlign with Tülu3-8B-SFT against four baselines: $\blacktriangle$ majority vote (MV) Tülu3-8B-DPO, and applied to Tülu3-8B-SFT the methods $\bullet$ best-of-$n$ (BoN), $\bullet$ MV, and $\bullet$ weighted MV (WMV) . All experiments use temperature 1.0 with reasoning included in model outputs. Note that Tülu3-8B-DPO model is the result of doing preference finetuning on the Tülu3-8B-SFT with 271k preference pairs. The costs associated with this process are not accounted for in this plot.[GF: change plot ]
Appendix H Reward Model Training Configuration
We trained a bt reward model using Llama-3.1-8B-Instruct (llama3) as the base model on a preference dataset for mathematical reasoning. The dataset contained 64 model-generated responses per prompt from GSM8K (gsm8k) (on-policy data from Llama-3.1-8B-Instruct), from which we constructed one preference pair per prompt by selecting one ground truth correct answer as preferred and one incorrect answer as dispreferred. Training hyperparameters are detailed in Table 2.
Table 2: Training Hyperparameters for the RM for the task-specific experiments.
Parameter Value
Model Configuration
Fine-tuning Type Full
Chat Template llama3
Dataset
Validation Split 5%
Optimization
Learning Rate $1\times 10^{-5}$
Total Batch Size 128
Training Epochs 1.0
LR Scheduler Linear Decay
Warmup Ratio 0.03
Weight Decay 0.0
Precision bfloat16
Infrastructure
Optimization Framework DeepSpeed ZeRO Stage 3
GPUS 8 L40S